OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6)

OpenStudy (anonymous):

8X-12+8X-32=4X+16 then combine like terms....

OpenStudy (anonymous):

HOW?

OpenStudy (anonymous):

you add 8X and 8X to get 16

OpenStudy (anonymous):

OOH OK gracias :)

OpenStudy (anonymous):

what me to tell you the answer?

OpenStudy (anonymous):

you want to give me the answer

OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6) 8X-12+8X-32=4X+12 16X-20=4X+12 -20 -20 16X= 4X-8 -4X _4X 10X= -8 divide by 10 on both sides X= -8/10 (reduce) X= -4/5

OpenStudy (anonymous):

thank u sahira

OpenStudy (anonymous):

Yeah no problem Roney! Hope I helped, because I know math is challenging...

OpenStudy (anonymous):

for real i need a lot help with math problems

OpenStudy (anonymous):

Yeah I get you, You're in Algebra 1 right? Just wait until you get to Algebra 2 (It's way harder). It's currently my 5th time taking Algebra 2

OpenStudy (anonymous):

im in algebra 2

OpenStudy (anonymous):

what the heck? That sure seems like Pre- Algebra or Algebra 1

OpenStudy (anonymous):

ooh hehe idk